# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ Abstract
This paper introduces and analyses interacting underdamped Langevin algorithms, termed Kinetic Interacting Particle Langevin Monte Carlo (KIPLMC) methods, for statistical inference in latent variable models. We propose a diffusion process that evolves jointly in the space of parameters and latent variables and exploit the fact that the stationary distribution of this diffusion concentrates around the maximum marginal likelihood estimate of the parameters. We then provide two explicit discretisations of this diffusion as practical algorithms to estimate parameters of statistical models. For each algorithm, we obtain nonasymptotic rates of convergence for the case where the joint log-likelihood is strongly concave with respect to latent variables and parameters. In particular, we provide convergence analysis for the diffusion together with the discretisation error, providing convergence rate estimates for the algorithms in Wasserstein-2 distance. We achieve accelerated convergence rates clearly demonstrating improvement in dimension dependence, similar to the underdamped samplers. To demonstrate the utility of the introduced methodology, we provide numerical experiments that demonstrate the effectiveness of the proposed diffusion for statistical inference and the stability of the numerical integrators utilised for discretisation. Our setting covers a broad number of applications, including unsupervised learning, statistical inference, and inverse problems.
@@ 1 Introduction
Parametric latent variable models (LVMs) are ubiquitous in many areas of statistical science, e.g., complex probabilistic models for text, audio, video and images [9, 65, 40] or inverse problems [66]. Intuitively, these models capture the underlying structure of the real data in terms of low-dimensional latent variables - a structure that often exists in real world [69]. However, latent variable models often require nontrivial procedures for maximum likelihood estimation as this quantity is often intractable [27]. Our main aim in this paper is to propose a set of accelerated and stable algorithms for the problem of maximum marginal likelihood estimation in the presence of latent variables and show convergence bounds for them by studying their nonasymptotic behaviour.
Let us consider a generic latent variable model as $p_{\theta}(x,y)$, parameterised by $\theta\in\mathbb{R}^{d_{\theta}}$, for fixed data $y\in{\mathbb{R}}^{d_{y}}$, and latent variables $x\in{\mathbb{R}}^{d_{x}}$. This model is considered for fixed observed data $y$, thus formally we see the statistical model as a real-valued mapping $p_{\theta}(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{\theta}}\to{\mathbb{R}}$. The task we are interested in is to estimate the parameter $\theta$ that explains the fixed dataset $y$. Often, this task is achieved via the maximum likelihood estimation (MLE). In accordance with this, in our setting, due to the presence of latent variables, we aim at finding the maximum marginal likelihood estimate, which termed as maximum marginal likelihood estimation (MMLE) problem [27]. More precisely, our problem takes the form
@@
where $p_{\theta}(y):=\int p_{\theta}(x,y)\mathrm{d}x$ is the marginal likelihood (also called the model evidence in Bayesian statistics [7]). It is apparent from (1) that the problem cannot be solved via optimisation techniques alone, as the marginal likelihood contains an intractable integral for most statistical models.
@@
Classically, this problem is solved with the iterative expectation maximisation (EM) algorithm [27], which provably converges to a local maximum. This algorithm consists of the E and M steps: (1) the E-step (optimally) produces an estimate for the posterior distribution of latent variables $p_{\theta}(x|y)$ and (2) the M-step maximises the expected log-likelihood estimate w.r.t. the parameter $\theta$. More compactly, starting at $\theta_{0}$, the EM algorithm produces the estimates $(\theta_{n})_{n\in{\mathbb{N}}}$, solving $\theta_{n+1}\in\operatornamewithlimits{argmax}_{\theta\in{\mathbb{R}}^{d_{\theta}}}\mathbb{E}_{p_{\theta_{n}}(x|y)}[\log p_{\theta}(x,y)]$. It is easy to establish that $\log p_{\theta_{n+1}}(y)\geq\log p_{\theta_{n}}(y)$, hence the method converges to a local maximum. In this setting, two tasks need to be solved simultaneously: (i) inference, i.e., inferring the distribution of $x$ given $y$ for $\theta_{n}$, in other words, the posterior distribution $p_{\theta_{n}}(x|y)$, (ii) estimation, i.e., estimating the maximiser $\theta_{n+1}$, whose computation requires the inference step. Often, this recursion is impossible to realise, due to two main challenges, namely, the expectation (E) and maximisation (M) steps, which are intractable. These steps can be approximated with a variety of methods, which, due to the simplicity of this algorithm, have been extensively explored. For example, Monte-Carlo EM [68] and stochastic EM [15] have been widely studied, see, e.g., [16, 17, 64, 10, 13, 28].
@@
In the most general case, the EM algorithm is implemented using Markov chain Monte Carlo (MCMC) techniques for the E-step, and numerical optimisation techniques for the M-step, e.g., [52, 50, 48]. With the popularity of unadjusted Langevin algorithm (ULA) in Bayesian statistics and machine learning [31, 33, 32, 22, 23, 25], variants of the implementation of expectation-maximisation algorithms using unadjusted chains to perform the E-step have also been proposed. Most notably, [26] studied an algorithm termed stochastic optimisation via unadjusted Langevin (SOUL), which performs unadjusted Langevin steps for the E-step and stochastic gradient ascent for the M-step, building on the ideas of [6]. This algorithm required to run a Markov chain for each E-step, resulting in a double-loop algorithm. The bias incurred by unadjusted chains complicates the theoretical analysis and requires a delicate balance of the step-sizes of the Langevin algorithm and gradient step to guarantee convergence [26].
@@
An alternative approach was developed in [47], where instead of running Markov chains to perform E-step, the authors proposed to use an interacting particle system, consisting of $N$ particles. Particle systems, such as the one proposed in [47], are often used to accelerate convergence for gradient-based methods or to replace these entirely ([11, 56, 46, 2, 39] for a review on their advantages and behaviour). In a similar vein, [47] shows that using an algorithm based on an interacting particle system (IPS), termed particle gradient descent (PGD), alleviates the need of running interleaving Markov chains, which can also lead to provable guarantees, see, e.g., [14]. Inspired by the approach in [47], a closely related interacting particle system was proposed in [3], where a properly scaled noise is injected in the $\theta$-dimension. This seemingly small modification is significant, making the algorithm an instance of a Langevin diffusion (an observation we build on in this paper). The authors termed this method interacting particle Langevin algorithm (IPLA) and proved error bounds for the algorithm. In particular, they showed that the parameter-marginal of IPLA iterates, targets a measure of the form $\pi_{\Theta}\propto\exp(-N\log p_{\theta}(y))$, which concentrates around the MMLE estimate with a rate $\mathcal{O}(N^{-1/2})$ where $N$ is the number of particles. The authors of [3] further provided the exponential convergence rate and the discretisation error under the within strongly convex setting. The IPLA methodology has been already utilised in other contexts, see, e.g., [45] for superlinear extensions and [35] for a set of proximal methods based on IPLA, and [4] for relations between IPLA and multiscale methods.
Finally, targeting this sort of measure to solve optimisation problems, is similar to simulated annealing techniques from optimisation (see [37, 42, 30, 43] for a discussion of these methods), which is a commonly proposed approach to MMLE. In particular, [30, 43] consider particle systems and observe sampling improvements in non-convex optimisation, motivating this alternative approach.
Contributions. In this paper, the overdamped Langevin diffusion proposed by [3] is modified by considering the underdamped setting - resulting in another stochastic differential equation (SDE) termed, kinetic interacting particle Langevin diffusion (KIPLD). In continuous time, our proposed diffusion corresponds to the noisy version of the system proposed in [49], and an acceleration of the overdamped diffusion introduced by [3], see, e.g., [51]. We then propose two underdamped Langevin samplers, which we term kinetic interacting particle Langevin Monte Carlo (KIPLMC), based on this diffusion. Our contributions can be summarised as follows:
@@
  • We propose the KIPLD, a diffusion process to optimise the marginal likelihood in latent variable models. Our diffusion process corresponds to $\theta$-noised version of the SDE proposed in [49]. However, in [49], the authors left the problem of nonasymptotic analysis open. In this paper, we provide a nonasymptotic analysis and show that the proposed modification enables us to prove that the stationary measure of this diffusion concentrates around the MMLE (Propositions 1 and 2) and the KIPLD converges to this measure exponentially fast (Proposition 3).
@@
  • We propose two discretisations of this diffusion to obtain practical algorithms, which we term KIPLMC methods. Our first algorithm KIPLMC1 corresponds to an exponential integrator discretisation of our SDE. This algorithm is also similar to the momentum particle gradient descent (MPGD) algorithm provided by [49] which does not have noise in $\theta$-dimension [^1] While [49] uses a gradient correction mechanism to stabilise the discretisation of the SDE, we directly analyse the discretisation without any gradient correction steps.. However, we provide discretisation error bounds for our method showing convergence rates in both time and step-size (Theorem 1), closing the problem of nonasymptotic analysis of this method in the strongly convex case. In particular, we show that KIPLMC1 attains accelerated rates of convergence compared to its overdamped counterparts IPLA [3] and PGD [47]. While these overdamped diffusion-based MMLE methods achieve an $\varepsilon$ error in Wasserstein-2 distance in $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps, we prove that KIPLMC1 accelerates this process, reaching $\varepsilon$ error in $\widetilde{\mathcal{O}}(\sqrt{d_{x}}\varepsilon^{-1})$ steps, similar to the performance of underdamped Langevin-based sampling algorithms [20, 25].
  • We then introduce a splitting-based explicit discretisation scheme within this setting, which leads to a nontrivial algorithm, which we term KIPLMC2. We provide a nonasymptotic analysis of this method (Theorem 2). We show that while MPGD algorithm provided by [49] only achieves stability with gradient correction steps, our KIPLMC2 discretisation attains similar stability while having theoretical guarantees.
  • Finally, we provide numerical results that show, in particular, the stability attained by KIPLMC2. We show in particular that this discretisation (which is based on [53]) is more stable than KIPLMC1 and MPGD of [49] when the latter is implemented as a standard discretisation of the diffusion provided in [49] without the use of the numerical trick called gradient correction. We show further that KIPLMC2 produces similar stable behaviour as the MPGD with gradient correction while also having solid theoretical guarantees.
@@
The paper is structured as follows. In Section 2, we present the ULA, its underdamped counterpart, and interacting particle systems for the MMLE problem. Following this introduction, in Section 3, we present a new diffusion targeting the solution of the MMLE problem and present the associated algorithms KIPLMC1 and KIPLMC2. Section 4 presents the nonasymptotic analysis of KIPLMC methods, providing a clear evidence of improved theoretical bounds due to the acceleration. Finally, we provide solid empirical evidence in the experiments of Section 5, showing that KIPLMC methods constitute a set of accelerated and theoretically backed tools to solve challenging statistical inference problems.
@@
[…]
Recall that, under a fixed dataset, our model is parameterised over $\theta\in\mathbb{R}^{d_{\theta}}$ and $x\in\mathbb{R}^{d_{x}}$. Let us consider an unnormalised probability model $U(\theta,x)$, where
$U(\theta,x)=-\log p_{\theta}(x,y).$
For any $p>0$ define the Wasserstein-$p$ metric as
$W_{p}(\pi,\nu)=\inf_{\Gamma\in\mathbf{T}(\pi,\nu)}\left(\int\|x-y\|^{p}_{p}\mathrm{d}\Gamma(x,y)\right)^{\frac{1}{p}},$
where $\mathbf{T}(\pi,\nu)$ denotes the set of couplings over $\mathbb{R}^{d\times d}$, with marginals $\pi$ and $\nu$.
@@ 2 Technical Background
Before presenting our proposed diffusion and algorithms, we introduce some concepts that will be useful to us in this paper.
2.1 Unadjusted Langevin Algorithm
Consider the problem of simulating random variables with a law in $\mathcal{P}({\mathbb{R}}^{d})$ of the form
$\pi({\mathrm{d}}z)\propto e^{-U(z)}{\mathrm{d}}z,$ (2)
where $U:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a potential function. This is a classical problem in computational statistics literature, with abundance of MCMC methods available, see, e.g., [58] for a book long treatment. In this paper, we focus on a particular class of MCMC algorithm, via the use of a discretisation of a particular SDE. The methods we are interested in rely on the fact that the following Langevin diffusion process
@@ 2.1 Unadjusted Langevin Algorithm
where $({\mathbf{B}}_{t})_{t\geq 0}$ is a standard Brownian motion, leaves the target measure (2) invariant. While this idea is well-known in MCMC literature and exploited for a long time as the proposal of the Metropolis adjusted Langevin algorithm (MALA) [60, 59, 70], recently several approaches simply drop the Metropolis step in this method and use the plain discretisation of the diffusion (3) as a sampling algorithm. This results in the following discretisation (termed unadjusted Langevin algorithm (ULA) [32, 31] or Langevin Monte Carlo (LMC) [22])
$Z_{n+1}^{\eta}=Z^{\eta}-\eta\nabla U(Z^{\eta}_{n})+\sqrt{2\eta}W_{n+1},$
where $(W_{n})_{n\in{\mathbb{N}}}$ are standard $d$-dimensional Normal random variables and $\eta>0$ corresponds to the step-size of the algorithm.
The performance and complexity of sampling with ULA, for convex and strongly convex $U$, have been extensively studied [31, 25]. The ULA has proved to be an efficient way to sample from these distributions, alleviating numerical and computational issues of the accept-reject steps. This improved computational performance comes at the expense of incurring a bias of order $\eta^{1/2}$ in the stationary measure of the algorithm [32]. Despite this, the ULA is a natural choice for sampling, as it can be implemented through a variety of discretisations and has, as discussed, favourable theoretical properties, see, e.g., [31, 67, 33, 12, 21, 60, 59] and [24], for convergence results under different assumptions on $U$.
2.2 Kinetic Langevin Monte Carlo
An alternative to the Langevin diffusions given in (3), which are generally termed overdamped Langevin diffusions, is another class of Langevin diffusions called underdamped diffusions [55]. This class of diffusions is akin to second-order differential equations and defined over position and momentum variables. The momentum dynamics are given by Newton’s Law for the motion of a particle, subject to friction and a stochastic forcing, acting as the gradient of the position particle [20]. In particular, for a $d$-dimensional target measure, the underdamped Langevin diffusion is defined to evolve on ${\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$ given by the SDE
@@ 2.2 Kinetic Langevin Monte Carlo
where $\gamma>0$ is called the friction coefficient and $({\mathbf{B}}_{t})_{t\geq 0}$ is a Brownian motion. From this equation, one can recover the overdamped dynamics in (3) by letting $\gamma\to\infty$ and considering the behaviour of the $z$-marginal [54, pg. 58]. Under certain regularity conditions [55], this system is known to be invariant w.r.t. an extended stationary measure of the form
@@
This means that we can recover the samples from our target measure (2) by sampling from (5) where the $z$-marginal of $\bar{\pi}$ is the target measure we would like to sample from.
Discretisation of (4) provides a set of algorithms termed as kinetic Langevin Monte Carlo (KLMC) [25]. These methods define discretisation methods tailored to the structure of the SDE given in (4), although Euler-Maruyama methods for this class of SDEs are also analysed [1, 38, 18]. As the update rules of the resulting algorithm are determined by a numerical approximation of a continuous-time process, how this algorithm performs is dependent on the smoothness of the sample paths and the mixing rate of the system. It is quite easy to see that the sample paths in the underdamped case are smoother, having an Hölder continuity of order $1+\alpha$, for $\alpha\in[0,1/2)$ (see introduction of [25]). Furthermore, similar to the ULA algorithm, this system also has exponentially fast convergence to the stationary measure [25]. By using an underdamped, second-order system, [20] exploit these properties to accelerate the first-order process, connecting this work to Nesterov acceleration and the growing body of literature dedicated to it. Indeed, [20] show that for certain choices of $\gamma$ the underdamped system converges faster to the stationary measure than the overdamped system and identify the regimes in which the underdamped algorithm outperforms the overdamped for their sampling schemes. This acceleration has not only been noted in $W_{2}$, by [25, 19] and others, but also in Kullback-Leibler (KL) divergence by [51]. The underdamped process leads to some interesting properties, such as an “induced regularisation” via damping effects, as well as, convergence with better dependence on dimension and step-size, as shown in [19]. Hence, the underdamped Langevin diffusion, presents a worthwhile starting point for sampling algorithms and represents a competitive alternative to the popular ULA.

### v3

@@ Abstract
This paper introduces and analyses interacting underdamped Langevin algorithms, termed Kinetic Interacting Particle Langevin Monte Carlo (KIPLMC) methods, for statistical inference in latent variable models. We propose a diffusion process that evolves jointly in the space of parameters and latent variables and show that the stationary distribution of this diffusion concentrates around the maximum marginal likelihood estimate of the parameters. We then provide two explicit discretisations of this diffusion as practical algorithms to estimate parameters of statistical models. For each algorithm, we obtain nonasymptotic rates of convergence in Wasserstein-2 distance for the case where the joint log-likelihood is strongly concave with respect to latent variables and parameters. We achieve accelerated convergence rates clearly demonstrating improvement in dimension dependence. To demonstrate the utility of the introduced methodology, we provide numerical experiments that illustrate the effectiveness of the proposed diffusion for statistical inference. Our setting covers a broad number of applications, including unsupervised learning, statistical inference, and inverse problems.
@@ 1 Introduction
Latent variable models (LVMs) are ubiquitous in many areas of statistical science, e.g., complex probabilistic models for text, audio, video and images [9, 49, 29], or inverse problems [28]. These models capture the underlying structure of the data in terms of low-dimensional variables – a structure that often exists in real world [52]. However, LVMs often require non-trivial procedures for their estimation as their likelihood is often intractable.
Consider a generic latent variable model as $p_{\theta}(x,y)$, parameterised by $\theta\in\mathbb{R}^{d_{\theta}}$, for fixed, observed data $y\in{\mathbb{R}}^{d_{y}}$, and latent variables $x\in{\mathbb{R}}^{d_{x}}$. Thus, formally, we see the statistical model as a real-valued mapping $p_{\theta}(x,y):\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{\theta}}\to{\mathbb{R}}$. The task we are interested in is to estimate the parameter $\theta$ that explains the fixed dataset $y$. Often, this is achieved via the maximum likelihood estimation (MLE). In our setting, due to the presence of latent variables, we aim at finding the maximum marginal likelihood estimation (MMLE), which is termed the MMLE problem [22]. More precisely, our problem takes the form
@@
where $p_{\theta}(y):=\int p_{\theta}(x,y)\mathrm{d}x$ is the marginal likelihood (also called the model evidence in Bayesian statistics [7]). It is apparent from (1) that the problem cannot be solved via optimisation techniques alone for most statistical models.
@@
Classically, this problem is solved with the iterative expectation maximisation (EM) algorithm [22], which provably converges to a local maximum. At a given parameter estimate $\theta_{n}$, the EM algorithm implements a map $\theta_{n}\mapsto\operatornamewithlimits{argmax}_{\theta\in{\mathbb{R}}^{d_{\theta}}}\mathbb{E}_{p_{\theta_{n}}(x|y)}[\log p_{\theta}(x,y)]$, which is guaranteed to increase the marginal likelihood at each iteration. This procedure requires an “E-step”, namely the computation (or estimate) of the expectation w.r.t. the posterior distribution of the latent variables $p_{\theta_{n}}(x|y)$, and an “M-step”, which is the maximisation of the expected log-likelihood w.r.t. $\theta$. These steps are intractable in general and can be approximated with a variety of methods, which have been extensively explored. For example, Monte Carlo EM [50] and stochastic EM [14] have been widely studied [13, 15, 48, 10, 11, 23].
@@
In the most general case, the EM algorithm is implemented using Markov chain Monte Carlo (MCMC) techniques for the E-step [6], and numerical optimisation techniques for the M-step [40, 38, 36]. With the popularity of unadjusted Langevin algorithm (ULA) in Bayesian statistics and machine learning [24, 25, 20], novel variants of EM algorithms use unadjusted chains. Most notably, De Bortoli et al. [21] studied an algorithm termed Stochastic Optimisation via Unadjusted Langevin (SOUL), which performs unadjusted Langevin steps for the E-step and stochastic gradient ascent for the M-step, building on the ideas of Atchadé et al. [6]. In the MCMC based EM setting, the bias incurred finite MCMC steps complicates the theoretical analysis and requires stringent conditions for step-sizes.
@@
An alternative approach was developed in Kuntz et al. [35] where the authors proposed an interacting particle system, consisting of $N$ particles to replace sequential MCMC steps. This framework, termed particle gradient descent (PGD), is based on the idea of joint optimisation and sampling to avoid a double-loop structure as in MCMC-based approaches, which results in an efficient algorithmic framework as well as one that is amenable to theoretical analysis Caprio et al. [12]. Inspired by the approach in Kuntz et al. [35], a closely related interacting particle system was proposed in Akyildiz et al. [1], where a scaled noise is injected in the $\theta$-dimension. The authors termed this method interacting particle Langevin algorithm (IPLA) and proved error bounds for the algorithm. This seemingly small modification is significant, making the algorithm an instance of a Langevin diffusion (an observation we build on in this paper), streamlining the theoretical analysis and opening the door to a wide variety of extensions with theoretical guarantees. The IPLA methodology has been already utilised in other contexts, see, Johnston et al. [33] for superlinear extensions and Cordero-Encinar et al. [17] for a set of proximal methods based on IPLA, and Akyildiz et al. [2] for relations between IPLA and multiscale methods.
Contributions. In this paper, we develop optimisation methods for MMLE based on underdamped Langevin samplers. Our contributions can be summarised as follows:
@@
(C1) We propose the kinetic interacting particle Langevin diffusion (KIPLD), a diffusion process to optimise the marginal likelihood in latent variable models. We prove that the stationary measure of this diffusion concentrates around the MMLE (Propositions 1 and 2) and the KIPLD converges to this measure exponentially fast (Proposition 3).
@@
(C2) We then develop our first kinetic interacting particle Langevin Monte Carlo (KIPLMC) method, termed KIPLMC1– an exponential integrator discretisation of KIPLD. We prove discretisation error bounds for our method showing convergence rates in both time and step-size (Theorem 1), closing the problem of nonasymptotic analysis of this method in the strongly convex case. In particular, we show that KIPLMC1 attains accelerated rates of convergence compared to its overdamped counterparts IPLA [1] and PGD [12]. While these overdamped diffusion-based MMLE methods achieve an $\varepsilon$ error in Wasserstein-2 distance in $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps, we prove that KIPLMC1 attains $\varepsilon$ error in $\widetilde{\mathcal{O}}(\sqrt{d_{x}}\varepsilon^{-1})$ steps.
(C3) We then introduce a splitting-based explicit discretisation scheme within this setting, which leads to a novel MMLE algorithm, termed KIPLMC2. We study the nonasymptotic behaviour of this method (Theorem 2), showing that the algorithm obtains an error $\varepsilon$ in Wasserstein-2 distance in $\widetilde{\mathcal{O}}(\sqrt{d_{x}d_{\theta}}\varepsilon^{-2})$ steps.
(C4) Finally, we provide solid empirical evidence of the performance of our methods in a variety of settings, including synthetic and realistic models. In the convex setting, we validate the $\mathcal{O}(N^{-1/2})$ concentration onto the maximiser $\bar{\theta}_{\star}$ and demonstrate the increased stability of the momentum-based methods. In the non-convex setting, we show the acceleration of the proposed methods, as well as the importance of the choice of numerical integrator, validating our methodological contribution.
Relations to existing work. The KIPLD is similar to the continuous time system on which the momentum particle gradient descent (MPGD) algorithm by Lim et al. [37] is based (see Eq. (21) in Lim et al. [37] where we take $\eta_{\theta}=\eta_{x}=1$ and $\gamma_{x}=\gamma_{\theta}=\gamma$). In line with the work done by Akyildiz et al. [1], we inject the $\theta$-dynamics with an appropriately scaled Brownian motion, i.e., $\sqrt{2\gamma/N}{\mathrm{d}}{\mathbf{B}}^{0}_{t}$. The advantage of studying and implementing this version is due to the fact that the new system can be shown to be an example of a standard underdamped Langevin diffusion, which enables us to show non-asymptotic bounds in a streamlined manner. Furthermore, our discretisations differ from the MPGD routine: the KIPLMC1 algorithm does not use a gradient correction step, whilst KIPLMC2 is based on a different integrator. In Sec. 5, we show that these algorithms obtain similar stability, while having theoretical guarantees.
@@
The paper is structured as follows. In Section 2, we present the background work on underdamped Langevin diffusions for sampling and interacting particle systems for the MMLE problem. Following this introduction, in Section 3, we present a new diffusion targeting the solution of the MMLE problem and present the associated algorithms KIPLMC1 and KIPLMC2. Section 4 presents the nonasymptotic analysis of KIPLMC methods, providing a clear evidence of improved theoretical bounds due to the acceleration. Finally, in Section 5, we present a variety of experiments to support our theoretical findings.
@@
Let $\mathcal{P}(\mathbb{R}^{d})$, for $d\geq 1$, denote the set of probability measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-algebra. We write $\langle\cdot,\cdot\rangle$ and $\|\cdot\|$ for the Euclidean inner product and norm on $\mathbb{R}^{d}$. We also use $N]=\{1,\dots,N\}$ and $\mathbb{N}$} for the positive integers. For $p>0$, define the Wasserstein-$p$ distance by
$W_{p}(\pi,\nu)=\left(\inf_{\Gamma\in\mathbf{T}(\pi,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\hskip-6.0pt\|x-y\|_{p}^{p}\,\mathrm{d}\Gamma(x,y)\right)^{1/p}$
where $\mathbf{T}(\pi,\nu)$ is the set of couplings on $\mathbb{R}^{d}\times\mathbb{R}^{d}$.
We also define $\widetilde{\mathcal{O}}$ as the big $\mathcal{O}$ notation, up to logarithmic orders.
@@ 2 Technical Background
Before presenting our proposed diffusion and algorithms, we introduce some concepts that will be useful to us in this paper. We first introduce the kinetic Langevin diffusion and its associated algorithms, which are the building blocks of our proposed methods. Then, we present the IPLA algorithm, which is the closest method to our proposed algorithms and serves as a starting point for our developments.
2.1 Kinetic Langevin Monte Carlo
In recent years, the Langevin diffusion and its associated algorithms have been widely used in statistics and machine learning to sample from complex distributions. The overdamped Langevin diffusion is given by the stochastic differential equation (SDE)
@@ 2.1 Kinetic Langevin Monte Carlo
where $U:{\mathbb{R}}^{d}\to{\mathbb{R}}$ is a potential function and $({\mathbf{B}}_{t})_{t\geq 0}$ is a Brownian motion. Under certain regularity conditions, this system is known to be invariant w.r.t. the measure $\pi({\mathrm{d}}z)\propto\exp(-U(z)){\mathrm{d}}z$ [43]. Under the strongly convex and Lipschitz-gradient setting, a standard Euler-Maruyama discretisation of this diffusion, called ULA, is known to attain at most $\mathcal{O}(\varepsilon)$ error in Wasserstein-2 distance to the stationary measure in $\widetilde{\mathcal{O}}(d\varepsilon^{-2})$ steps [20, 25].
An alternative to the overdamped Langevin diffusions given in (2), is another class of diffusions called underdamped Langevin diffusions [43]. This class of diffusions is akin to second-order differential equations and defined over position and momentum variables. In particular, for a $d$-dimensional target measure $\pi({\mathrm{d}}z)\propto\exp(-U(z)){\mathrm{d}}z$, the underdamped (kinetic) Langevin diffusion is given by the SDE
@@
where $\gamma>0$ is called the friction coefficient and $({\mathbf{B}}_{t})_{t\geq 0}$ is a Brownian motion. Under certain regularity conditions [43], this system is known to be invariant w.r.t. an extended stationary measure of the form
@@
This means that we can recover the samples from our target measure $\pi({\mathrm{d}}z)\propto\exp(-U(z)){\mathrm{d}}z$ by sampling from (4) and then marginalising out the velocity variable.
The structure in (3) allows for smoother sample paths and faster convergence to the stationary measure compared to the overdamped case [16]. This has motivated the development of algorithms based on discretisations of the underdamped Langevin diffusion, termed kinetic Langevin Monte Carlo (KLMC) [19]. These methods have been shown to achieve improved convergence rates compared to their overdamped counterparts, making them a competitive alternative for sampling from complex distributions. In particular, under the strongly convex and Lipschitz-gradient setting, KLMC is known to attain at most $\mathcal{O}(\varepsilon)$ error in Wasserstein-2 distance to the stationary measure in $\widetilde{\mathcal{O}}(\sqrt{d}\varepsilon^{-1})$ steps [19]. In this paper, we build on the framework of KLMC methods to develop algorithms for the MMLE problem.